solve the following system of equations.
x+7y=9
x=4-7y

Answer by Alan3354(69443)

solve the following system of equations.
x+7y=9
x=4-7y --> x+7y = 4
----------------
x+7y = 9
x+7y = 4
--------------- Subtract
0 = 5
---
It's inconsistent ---> no solution
If you graph the equations, you get parallel lines.
No intersection, no solution